# Random coverage from within with variable radii, and Johnson-Mehl cover timesthanks: Supported by EPSRC grant EP/T028653/1

### v2

@@ 4 Proof of results for the J-M model
By our choice of $\rho$ and the Mapping theorem (see e.g. [6]), $L{\cal H}_{\rho}$ is a homogeneous Poisson process of intensity $1$ on $\mathbb{R}^{d}\times\mathbb{R}_{+}$, and therefore by the definition (2.10), $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$.
@@
$\lim_{\rho\to\infty}\left(\omega_{d}\rho t_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\right)=\beta,$ (4.1)
@@
and by (4.1), as $\rho\to\infty$ this tends to
@@
Hence we have the condition (3.2) from Lemma 3.2 (with $\beta$ on the right hand side replaced by the last displayed expression). Also $\mathbb{E}\,[Y^{d-1}]=1/d$ and $\mathbb{E}\,[Y^{d}]=1/(d+1)$.
@@
This yields (2.11).
For (2.12), take $\rho=L^{d+1}$, so that $\log\log L=\log\log\rho-\log(d+1)$. By Lemma 4.1, $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$. Therefore
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L\leq\beta\right]$
$=\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho+d^{2}\log(d+1)\leq\beta\right],$
and by (2.11), this tends to $\exp(-c_{d}(d^{d}\omega_{d})^{-1/(d+1)}(d+1)^{d^{2}/(d+1)}|A|e^{-\beta/(d+1)}),$ and then (2.12) follows. ∎
We now derive Theorems 2.9 and 2.10 from Theorems 2.1 and 2.2, respectively.
@@
and hence (2.13).
For (2.14), let $\rho=L^{3}$, so that $\log\log\rho=\log\log L+\log(3)$. Then by Lemma 4.1,
$\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L\leq\beta]=\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho+\log 81\leq\beta],$
and by (2.13), this converges to $\exp(-(4\pi)^{-1/3}|A|3^{4/3}e^{-\beta/3}-(2\pi^{2})^{-1/3}|\partial A|3^{4/6}e^{-\beta/6}),$ and hence (2.14). ∎
@@
$t_{\rho}:=\Big{(}\frac{(2(d-1)\log\rho+2d(d-1)\log\log\rho+\beta)\vee 0}{\omega_{d}\rho}\Big{)}^{1/(d+1)},$ (4.2)
@@
and $\log\log(\rho t_{\rho})=\log\log\rho+\log(d/(d+1))+o(1).$ Therefore taking $Y$ to be uniformly distributed on $[0,1]$, so that $\mathbb{E}\,[Y^{d}]=1/(d+1)$, using (4.2) we obtain that
@@
×$\times\Big{(}\frac{d}{d+1}\Big{)}^{d-2+1/d}e^{-\beta/(2d+2)}\Big{)}$

### v3

@@ 3 Proof of results for the J-M model
By our choice of $\rho$ and the Mapping theorem (see e.g. [9]), $L{\cal H}_{\rho}$ is a homogeneous Poisson process of intensity $1$ on $\mathbb{R}^{d}\times\mathbb{R}_{+}$, and therefore by the definition (5), $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$.
@@
$\lim_{\rho\to\infty}\left(\omega_{d}\rho t_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\right)=\beta,$ (21)
@@
and by (21), as $\rho\to\infty$ this tends to
@@
Hence we have the condition (19) from Lemma 3.1 (with $\beta$ on the right hand side replaced by the last displayed expression). Also $\mathbb{E}\,[Y^{d-1}]=1/d$ and $\mathbb{E}\,[Y^{d}]=1/(d+1)$.
@@
This yields (6).
For (7), take $\rho=L^{d+1}$, so that $\log\log L=\log\log\rho-\log(d+1)$. By Lemma 3.2, $L\tilde{T}_{\rho}$ has the same distribution as $\tilde{\tau}_{L}$. Therefore
$\mathbb{P}\left[\omega_{d}\tilde{\tau}_{L}^{d+1}-d(d+1)\log L-d^{2}\log\log L-d^{2}\log(d+1)\leq\beta\right]$
$=\mathbb{P}\left[\omega_{d}\rho\tilde{T}_{\rho}^{d+1}-d\log\rho-d^{2}\log\log\rho\leq\beta\right],$
and by (6), this gives us (7).
We now derive Theorems 2.2 and 2.3 from Theorems 2.9 and 2.10, respectively.
@@
and hence (8).
For (9), let $\rho=L^{3}$, so that $\log\log\rho=\log\log L+\log(3)$. Then by Lemma 3.2,
$\mathbb{P}[\pi\tau_{L}^{3}-6\log L-4\log\log L-\log 81\leq\beta]=\mathbb{P}[\pi\rho T_{\rho}^{3}-2\log\rho-4\log\log\rho\leq\beta],$
and by (8), this yields (9).
@@
$t_{\rho}:=\Big(\frac{(2(d-1)\log\rho+2d(d-1)\log\log\rho+\beta)\vee 0}{\omega_{d}\rho}\Big)^{1/(d+1)},$ (22)
@@
and $\log\log(\rho t_{\rho})=\log\log\rho+\log(d/(d+1))+o(1).$ Therefore taking $Y$ to be uniformly distributed on $[0,1]$, so that $\mathbb{E}\,[Y^{d}]=1/(d+1)$, using (22) we obtain that
@@
$\times\Big(\frac{d}{d+1}\Big)^{d-2+1/d}e^{-\beta/(2d+2)}\Big)$